OpenStudy (anonymous):

Part 1: Solve |2x-9|+4_> 7.(its the line under the sign) You must show all work. Use >= and <= to represent the inequalities. (3 points) Part 2: Use complete sentences to describe the graph in words. (3 points) Answer

OpenStudy (hoblos):

|2x-9| >= 3 => 2x-9 >= 3 2x >= 12 x>= 6 or 2x-9 <= -3 2x <= 6 x <= 3

OpenStudy (anonymous):

\[|2x - 9|+4\ge7\] \[|2x-9|\ge7-4\] \[2x \ge3+9\] \[x \ge12/2\] \[x \ge6\]